### v3

@@ 5.2 Labeling Algorithm for the ARMP Pricing Problem
[…]
@@
$T_{v^{\prime}}^{\mathrm{cost}}$ $=T_{v}^{\mathrm{cost}}+W_{v,v^{\prime}}(F_{v}^{L,\mathrm{max}})$
$T_{v^{\prime},i}^{\mathrm{perf}}$ $=\begin{cases}\begin{aligned} &T_{v,i}^{\mathrm{perf}}-1&&\text{if}\ \ i=v^{\prime}\\ &1&&\text{if}\ \ F_{v^{\prime}}^{L^{\prime},\mathrm{max}}+t_{v^{\prime},i}(\omega_{\mathrm{max}})>\mathrm{LF}_{i}^{\mathrm{e}}-p_{i,q}\\ &1&&\text{if}\ \ F_{v^{\prime}}^{L^{\prime},\mathrm{max}}+t_{v^{\prime},i}^{\alpha}>\mathrm{LF}_{i}-p_{i,q}\\ &T_{v,i}^{\mathrm{perf}}&&\text{otherwise}\end{aligned}\end{cases}$
$F_{v^{\prime}}^{L^{\prime}}(\tau)$ $=\begin{cases}\begin{aligned} &\underset{z=0}{\overset{\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}}{\sum}}\mathbb{P}(F_{v}^{L}+t_{v,v^{\prime}}+p_{v^{\prime},q}=z)&&\text{if}\ \ \tau=\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}\\ &\mathbb{P}(F_{v}^{L}+t_{v,v^{\prime}}+p_{v^{\prime},q}=\tau)&&\text{if}\ \ \tau>\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}\\ &0&&\text{otherwise}\end{aligned}\end{cases}$
$F_{v^{\prime}}^{L^{\prime},\mathrm{max}}$ $=\max\left\{F_{v}^{L,\mathrm{max}}+t_{v,v^{\prime}}(\omega_{\mathrm{max}}),\mathrm{ES}_{v^{\prime}}\right\}+p_{v^{\prime},q}$
@@
Definition 5.1 (Dominance Rule)
[…]
@@
[…]
@@
$F_{v}^{1,\mathrm{max}}$ $=F_{v}^{2,\mathrm{max}}$ (18)
Properties (16) and (17) ensure that any feasible extension of $L^{2}$ is also feasible for $L^{1}$. Constraints (15), (17) and (18) guarantee that, after extending both $L^{1}$ and $L^{2}$ along $(v,v^{\prime})$, the reduced cost of $L^{1^{\prime}}$ is still less or equal to the reduced cost of $L^{2^{\prime}}$.
[…]
$\left[\mathrm{ES}_{i}-t_{o,i}(\omega_{\mathrm{max}}),\mathrm{LF}_{i}-p_{i,q}-t_{o,i}^{\alpha}\right].$ (19)
Leaving the depot at a time instant greater than $\mathrm{LF}_{i}-p_{i,q}-t_{o,i}^{\alpha}$ would violate chance constraint (3) at task $i$, while leaving before $\mathrm{ES}_{i}-t_{o,i}(\omega_{\mathrm{max}})$ incurs unnecessary waiting time. These labels are also called initial labels. We then extend these labels using the previously described resources extension function, discard dominated labels, and repeat the same procedure for the oldest label until no feasible extensions can be found anymore. We note that the calculation of the start time distribution consumes the majority of runtime during label extensions. When the support of travel time distributions is small enough, these calculations can be done exactly. We refer to Errico et al. (2016) and Errico et al. (2018) for an extensive description of such an algorithm.
@@ 5.3 Peculiarities for the DRMP
Though the DRMP is structurally very similar to the ARMP, several characteristics of the former must be considered when solving the pricing problem of the DRMP. Because the node and arc set does not depend on the skill composition $s$, we can define the graph of a disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ as $\mathcal{G}^{q,s}\coloneqq(\mathcal{V}^{q},\mathcal{A}^{q},\beta)$, where the only difference lies in the coefficients $\beta_{k,t}^{s}(\omega_{\mathrm{max}})$ replacing $b_{k,t}^{q}(\omega_{\mathrm{max}})$ during the calculation of the time-dynamic arc weights. We note that $\beta\coloneqq\left(\beta_{k,t}^{s}(\omega_{\mathrm{max}})\right)_{k\in\mathcal{K},t\in\mathcal{T}}$ denotes the disaggregated workforce requirements of a team with profile $q$ and skill composition $s$.
@@
Definition 5.2 (Dominance rule for the DRMP)
@@
[…]
@@
$T_{v}^{1,\mathrm{cost}}+\underset{k\in\mathcal{K}}{\sum}\underset{\tau=\mathrm{tl}^{1}}{\overset{F_{v}^{1,\mathrm{max}}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{P^{1},s,1}(\omega_{\mathrm{max}})$ $\leq T_{v}^{2,\mathrm{cost}}+\underset{k\in\mathcal{K}}{\sum}\underset{\tau=\mathrm{tl}^{2}}{\overset{F_{v}^{2,\mathrm{max}}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{P^{2},s,2}(\omega_{\mathrm{max}})$ (20)
[…]
@@
$F_{v}^{1,\mathrm{max}}$ $=F_{v}^{2,\mathrm{max}}$ (23)
@@
Dominance rule 5.2 offsets the reduced costs of $L^{1}$ and $L^{2}$ by the workforce penalty of each path and imposes an additional restriction (24) on their depot leave times. It is easy to see that if $L^{1}$ dominates $L^{2}$ in $\mathcal{G}^{q,s}$, it also dominates $L^{2}$ with respect to the dominance rule described in Definition 5.1. However, the converse is not true.
Figure 2: Labels satisfying dominance rule 5.1 and violating rule 5.2
Example 5.3
Figure 2 visualizes the pricing network for a disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ and two labels $L^{1}$ (solid line) and $L^{2}$ (dotted line). We assume $\mathcal{K}=\{1\}$, $s_{q,1}=1$ and $\delta_{1,\tau}=-1$ for all $\tau\in[0,10]$. Furthermore, we simplify our considerations by neglecting time window restrictions and setting $w_{1}=w_{2}=1$. Moreover, we define $p_{2,q}=3$ and $p_{1,q}=3$. The reduced costs for labels $L^{1}$ and $L^{2}$ are then equal to $T_{1}^{1,\mathrm{cost}}=13.5$ and $T_{2}^{2,\mathrm{cost}}=18$. Therefore, it is clear that $L^{1}$ dominates $L^{2}$ in the sense of Definition 5.1. However, if we offset the reduced cost of each label by the workforce penalties, we obtain
$T_{1}^{2,\mathrm{cost}}+\underset{\tau=0}{\overset{10}{\sum}}\delta_{1,\tau}\beta_{k,\tau}^{P^{2},s,2}(\omega_{\mathrm{max}})=18-11=7<7.5=13.5-6=T_{1}^{1,\mathrm{cost}}+\underset{\tau=5}{\overset{10}{\sum}}\delta_{1,\tau}\beta_{k,\tau}^{P^{1},s,1}(\omega_{\mathrm{max}})$
Hence, $L^{1}$ does not dominate $L^{2}$ in $\mathcal{G}^{q,s}$, i.e., in the sense of Definition 5.2. The dominance rule presented here tends to be more strict, leading to around 25% less dominated labels.
[…]
that is, all properties besides the reduced cost $T_{v}^{\mathrm{cost}}$ of $L$ remain the same. As the dominance rule on $\mathcal{G}^{q,s}$ merely considers reduced costs after offsetting them by the workforce penalty, it is clear that if a label $L^{1}$ dominates another label $L^{2}$ in $\mathcal{G}^{q,s}$ if and only if it dominates $L^{2}$ in $\mathcal{G}^{q,\tilde{s}}$ for all $\tilde{s}\in\mathcal{S}_{q}$. For a formal proof of these relations, see Appendix 12.
@@
In the application at hand, the set of pricing networks is given by $\mathcal{N}\coloneqq\left\{\mathcal{G}^{q,s}:\ q\in\mathcal{Q},s\in\mathcal{S}^{q}\right\}$ and its partition is $\mathcal{N}_{q}\coloneqq\left\{\mathcal{G}^{q,s}:\ s\in\mathcal{S}^{q}\right\}$ for all $q\in\mathcal{Q}$. Furthermore, path feasibility is defined by the satisfaction of constraints (3)–(4), which do not depend on the skill composition, i.e., (i) is satisfied by the definition of $\mathcal{N}_{q}$. Moreover, it is clear to see that the adjusted dominance rule (20)–(24) does not depend on the skill composition, which is the only property that is not shared among all pricing networks within $\mathcal{N}_{q}$. Finally, the usage of offset reduced costs (20), jointly with inequality (24), allows for the proof of dominance with respect to the original dominance rule (15)–(18) in all pricing networks in $\mathcal{N}_{q}$ if dominance with respect to (20)–(24) is satisfied in an arbitrary pricing network in $\mathcal{N}_{q}$. For additional details on how the equivalence of pricing networks can be derived using properties i)–iii), we refer to the proof presented in Appendix 12.

### v4

@@ 5.2 Labeling Algorithm for the ARMP Pricing Problem
We solve the pricing problem using a labeling algorithm with a customized dominance rule, enhanced with several acceleration strategies that reduce the graph size and dimension of the label space. In order to check if a path $P$ in $\mathcal{G}^{q}$ violates constraints (3) or (4) at a task $i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}$, full information about the distribution of $F_{i}^{P}$ and $F_{i}^{P,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{\gamma}}}$ needs to be available. Thus, for each partial path $P=(o,v_{1},\dots,v_{l},v)$ in $\mathcal{G}^{q}$, we define a label $L$ as a tuple $L\coloneqq\left(\mathrm{tl}^{L},v,P,T_{v}^{\mathrm{cost}},(T_{v,i}^{\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{L}\right)$, where $\mathrm{tl}^{L}$ is the depot leave time, $T_{v}^{\mathrm{cost}}$ is the reduced cost of the path, $T_{v,i}^{\mathrm{perf}}\in\{0,1\}$ indicate if a task node $i$ can still be visited and $F_{v}^{L}$ is the distribution of finish times at the current last node $v$. We say that $L$ is feasible for $\mathcal{G}^{q}$ if its associated path $P=(o,v_{1},\dots,v_{l},v)$ satisfies constraints (3) and (4) for all task nodes visited by $P$ and $v_{i}\neq v_{j}$ holds for all $i,j\in\{1,\dots,l\},i\neq j$.
When extending a label $L$ along arcs $(v,v^{\prime})$, we obtain a new label $L^{\prime}=\left(\mathrm{tl}^{L^{\prime}},v^{\prime},P^{\prime},T_{v^{\prime}}^{\mathrm{cost}},(T_{v^{\prime},i}^{\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v^{\prime}}^{L^{\prime}}\right)$ by using the following resource extension function:
@@
$T_{v^{\prime}}^{\mathrm{cost}}$ $=T_{v}^{\mathrm{cost}}+W_{v,v^{\prime}}(F_{v}^{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})})$
$T_{v^{\prime},i}^{\mathrm{perf}}$ $=\begin{cases}\begin{aligned} &T_{v,i}^{\mathrm{perf}}-1&&\text{if}\ \ i=v^{\prime}\\ &1&&\text{if}\ \ F_{v^{\prime}}^{L^{\prime}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{max}})+t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}_{v^{\prime},i}(\omega_{\mathrm{max}})>\mathrm{LF}_{i}^{\mathrm{e}}-p_{i,q}\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall k=1\,\dots\mathcal{B}}\\ &1&&\text{if}\ \ F_{v^{\prime}}^{L^{\prime}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{max}})+t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}_{v^{\prime},i}(\omega_{\alpha})>\mathrm{LF}_{i}-p_{i,q}\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall k=1\,\dots\mathcal{B}}\\ &T_{v,i}^{\mathrm{perf}}&&\text{otherwise}\end{aligned}\end{cases}$
$F_{v^{\prime}}^{L^{\prime}}(\tau)$ $=\begin{cases}\begin{aligned} &\underset{z=0}{\overset{\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}}{\sum}}\mathbb{P}(F_{v}^{L}+t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}^{L,\gamma}}}_{v,v^{\prime}}+p_{v^{\prime},q}=z)&&\text{if}\ \ \tau=\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}\\ &\mathbb{P}(F_{v}^{L}+t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}^{L,\gamma}}}_{v,v^{\prime}}+p_{v^{\prime},q}=\tau)&&\text{if}\ \ \tau>\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}\\ &0&&\text{otherwise}\end{aligned}\end{cases}$
@@
Example 2.
Consider the graph $\mathcal{G}^{q}$ for a given profile $q\in\mathcal{Q}$ depicted in Figure 2. Let $L^{1}$ be a label describing the subpath $(o,1)$ with depot leave time $tl^{1}=2$. We assume $\mathcal{K}=\{1\}$, $\xi_{q,1}=1$ and $\delta_{1,\tau}=-1$ for all $\tau\in[2,10]$. Furthermore, assume $LF^{e}_{1}=10$, $ES_{1}=0$, $ES_{3}=3$, $LF_{3}=LF^{3}_{e}=15$, $w_{1}=w_{3}=1$, $\mu_{1}=2$, and $\mu_{1}=2$. Moreover, we define $p_{1,q}=p_{3,q}=3$. Finally, define $\gamma=0.95$ and time bins $\mathcal{B}=\{\mathcal{B}_{1},\mathcal{B}_{2}\}$ with $\mathcal{B}_{1}=\{0,\dots 5\}$ and $\mathcal{B}_{2}=\{6,\dots,11\}$.
Figure 2: Label extension and time-dynamic arc weights
Because $tl^{1}=2$ and $p_{1,q}=3$ hold, the finish time distribution $F_{1}^{1}$ is then equal to $P(F_{1}^{1}=6)=P(F_{1}^{1}=7)=\dfrac{1}{2}$. Hence, the median finish time is $\bar{F_{1}^{1}}=6$. Furthermore, in scenario $\omega_{\gamma}$, we have $t^{1}_{o,1}(\omega_{\gamma})=2$. Thus, one can conclude that the finish time of $L^{1}$ at task $1$ in scenario $\omega_{\gamma}$ is equal to $F_{1}^{1}(\omega_{\gamma})=\max\{tl^{1}+t_{o,1}^{1}(\omega_{\gamma}),ES_{1}\}+p_{1,q}=2+2+3=7$. The reduced cost of label $L^{1}$ is then equal to
$T_{1}^{cost}=\mathbb{E}(F_{1}^{1})-\mu_{1}-\underset{t=2}{\overset{7}{\sum}}\delta_{1,\tau}=\dfrac{1}{2}\cdot(6+7)-2-\underset{t=2}{\overset{7}{\sum}}\delta_{1,t}=4.5+6=10.5$
Assume one wants to extend $L^{1}$ along the edge $(1,3)$. Because $\bar{F}_{1}^{1}=6\in\mathcal{B}_{2}$ holds, the finish time distribution $F_{1}^{3}$ thus depends on the travel time distribution of edge $(1,3)$ in time bin $\mathcal{B}_{2}$, i.e., on the distribution of $t_{1,3}^{2}$. Furthermore, the time-dynamic arc weight $W_{1,3}\left(F_{1}^{1}(\omega_{\gamma})\right)$ also depends on the distribution of $t_{1,3}^{2}$, as well as on the finish time distribution $F_{1}^{1}$. The finish time distribution at task $3$ is then given by $\mathbb{P}(F_{1}^{3}=10)=\mathbb{P}(F_{1}^{3}=11)=\dfrac{1}{2}$. Hence, the reduced cost of the resulting label is equal to
$T_{3}^{cost}=T_{1}^{cost}+W_{1,3}\left(F_{1}^{1}(\omega_{\gamma})\right)=10.5+\dfrac{1}{2}(10+11)-\underset{t=8}{\overset{10}{\sum}}\delta_{1,t}=10.5+10.5+3=24$
Moreover, task resources $T_{3,3}^{perf}$ and $T_{3,1}^{perf}$ are updated to values 0 and $1$, respectively. While the former is an immediate result of the extension along edge $(1,3)$, the latter can be reset to $1$ because
$F_{3}^{1}(\omega_{max})+t^{2}_{3,1}(\omega_{max})=11+t^{2}_{3,1}(\omega_{max})>12=15-3=LF_{1}^{e}-p_{1,q}$
holds for any non-negative travel time $t^{2}_{3,1}(\omega_{max})$ and thus, condition 2 of the task resource update in the resource extension function is satisfied. All remaining resources can be updated canonically according to the resource extension function.
@@
Definition 1 (Dominance Rule).
Let $L^{1}\coloneqq\left(\mathrm{tl}^{1},v,P^{1},T_{v}^{1,\mathrm{cost}},(T_{v,i}^{1,\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{1}\right)$ and
$L^{2}\coloneqq\left(\mathrm{tl}^{2},v,P^{2},T_{v}^{2,\mathrm{cost}},(T_{v,i}^{2,\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{2}\right)$ be labels in $\mathcal{G}^{q}$. We say that $L^{1}$ dominates $L^{2}$ if the following properties hold:
@@
$T_{v,i}^{1,\mathrm{perf}}$ $\geq T_{v,i}^{2,\mathrm{perf}}\quad\forall i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}$ (17)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{v}^{1}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{v}^{2}}$ (19)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{v}^{1}(\omega_{\gamma})}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{v}^{2}(\omega_{\gamma})}$ (20)
where $\bar{F}_{v}^{1}$ and $\bar{F}_{v}^{2}$ are the median finish times of labels $L^{1}$ and $L_{2}$ at task $v$, respectively.
Properties (17) and (18) ensure that any feasible extension of $L^{2}$ is also feasible for $L^{1}$. Condition (19) ensures that stochastic dominance (18) propagates along label extensions. Jointly with constraints (16), (18) and (20), this guarantees that, after extending both $L^{1}$ and $L^{2}$ along $(v,v^{\prime})$, the reduced cost of $L^{1^{\prime}}$ is still less or equal to the reduced cost of $L^{2^{\prime}}$.
At the beginning of each iteration of the labeling algorithm, we create one label for each task $i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}$ and each depot leave time for which the corresponding label satisfies (3)–(4). These labels are also called initial labels. We then extend these labels using the previously described resources extension function, discard dominated labels, and repeat the same procedure for the oldest label until no feasible extensions can be found anymore. We note that the calculation of the start time distribution consumes the majority of runtime during label extensions. When the support of travel time distributions is small enough, these calculations can be done exactly. We refer to errico_recourse and errico2018vehicle for an extensive description of such an algorithm.
@@ 5.3 Peculiarities for the DRMP
Though the DRMP is structurally very similar to the ARMP, several characteristics of the former must be considered when solving the pricing problem of the DRMP. Because the node and arc set does not depend on the skill composition $s$, we can define the graph of a disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ as $\mathcal{G}^{q,s}\coloneqq(\mathcal{V}^{q},\mathcal{A}^{q},\beta)$, where the only difference lies in the coefficients $\beta_{k,t}^{s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}}))$ replacing $b_{k,t}^{q}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})$ during the calculation of the time-dynamic arc weights. We note that $\beta\coloneqq\left(\beta_{k,t}^{s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})\right)_{k\in\mathcal{K},t\in\mathcal{T}}$ denotes the disaggregated workforce requirements of a team with profile $q$ and skill composition $s$.
@@
Definition 2 (Dominance rule for the DRMP).
@@
$L^{k}\coloneqq\left(\mathrm{tl}^{k},v,P^{k},T_{v}^{k,\mathrm{cost}},(T_{v,i}^{k,\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{k}\right)$
@@
$T_{v}^{1,\mathrm{cost}}+\underset{k\in\mathcal{K}}{\sum}\underset{\tau=\mathrm{tl}^{1}}{\overset{F_{v}^{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{P^{1},s,1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})$ $\leq T_{v}^{2,\mathrm{cost}}+\underset{k\in\mathcal{K}}{\sum}\underset{\tau=\mathrm{tl}^{2}}{\overset{F_{v}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{P^{2},s,2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})$ (21)
$T_{v,i}^{1,\mathrm{perf}}$ $\geq T_{v,i}^{2,\mathrm{perf}}\quad\forall i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}$ (22)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{v}^{1}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{F}_{v}^{2}}$ (24)
$F_{v}^{1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})$ $=F_{v}^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega_{\gamma}})$ (25)
@@
where $\bar{F}_{v}^{1}$ and $\bar{F}_{v}^{2}$ are the median finish times of labels $L^{1}$ and $L_{2}$ at task $v$, respectively.
Dominance rule 2 offsets the reduced costs of $L^{1}$ and $L^{2}$ by the workforce penalty of each path and imposes an additional restriction (26) on their depot leave times. It is easy to see that if $L^{1}$ dominates $L^{2}$ in $\mathcal{G}^{q,s}$, it also dominates $L^{2}$ with respect to the dominance rule described in Definition 1. However, the converse is not true.
Figure 3: Labels satisfying dominance rule 1 and violating rule 2
Example 3.
Figure 3 visualizes the pricing network of Figure 2 alongside an additional path $L^{2}$. Its path is denoted by dotted lines and equals the subtour $(o,2,3)$. In addition to the assumptions made in Example 2, we set $p_{2,q}=3$, $tl^{2}=0$, $\mu_{2}=0$, $\delta_{1,\tau}=-2$ for $\tau=0,1$, $s_{q,1}=1$, $w_{2}=1$, and $LF_{2}^{e}=10$. The reduced costs for labels $L^{1}$ and $L^{2}$ are then equal to $T_{1}^{1,\mathrm{cost}}=24$ and $T_{2}^{2,\mathrm{cost}}=27.5$. Therefore, it is clear that $L^{1}$ dominates $L^{2}$ in the sense of Definition 1. However, if we offset the reduced cost of each label by the workforce penalties, we obtain
$T_{1}^{2,\mathrm{cost}}+\underset{\tau=0}{\overset{10}{\sum}}\delta_{1,\tau}\beta_{k,\tau}^{P^{2},s,2}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})=27.5-13=14.5<15=24-9=T_{1}^{1,\mathrm{cost}}+\underset{\tau=2}{\overset{10}{\sum}}\delta_{1,\tau}\beta_{k,\tau}^{P^{1},s,1}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})$
Hence, $L^{1}$ does not dominate $L^{2}$ in $\mathcal{G}^{q,s}$, i.e., in the sense of Definition 2. The dominance rule presented here tends to be more strict, leading to around 25% less dominated labels.
As the only difference between the pricing graphs $\mathcal{G}^{q,\tilde{s}}$ and $\mathcal{G}^{q,s}$ for $q\in\mathcal{Q}$ and $s,\tilde{s}\in\mathcal{S}_{q}$ lies in the workforce penalty for the time-dynamic arc weights, we can transfer any feasible label $L=\left(\mathrm{tl}^{L},v,P,T_{v}^{\mathrm{cost}},(T_{v,i}^{\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{L}\right)$ in $\mathcal{G}^{q,s}$ to a different pricing network $\mathcal{G}^{q,\tilde{s}}$ and obtain a new feasible label $\tilde{L}$ in $\mathcal{G}^{q,\tilde{s}}$ with
$\tilde{L}=\left(\mathrm{tl}^{L},v,P,\tilde{T}_{v}^{\mathrm{cost}},(T_{v,i}^{\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{L}\right)$
that is, all properties besides the reduced cost $T_{v}^{\mathrm{cost}}$ of $L$ remain the same. As the dominance rule on $\mathcal{G}^{q,s}$ merely considers reduced costs after offsetting them by the workforce penalty, it is clear that if a label $L^{1}$ dominates another label $L^{2}$ in $\mathcal{G}^{q,s}$ if and only if it dominates $L^{2}$ in $\mathcal{G}^{q,\tilde{s}}$ for all $\tilde{s}\in\mathcal{S}_{q}$. For a formal proof of these relations, see E.
@@
In the application at hand, the set of pricing networks is given by $\mathcal{N}\coloneqq\left\{\mathcal{G}^{q,s}:\ q\in\mathcal{Q},s\in\mathcal{S}^{q}\right\}$ and its partition is $\mathcal{N}_{q}\coloneqq\left\{\mathcal{G}^{q,s}:\ s\in\mathcal{S}^{q}\right\}$ for all $q\in\mathcal{Q}$. Furthermore, path feasibility is defined by the satisfaction of constraints (3)–(4), which do not depend on the skill composition, i.e., (i) is satisfied by the definition of $\mathcal{N}_{q}$. Moreover, it is clear to see that the adjusted dominance rule (21)–(26) does not depend on the skill composition, which is the only property that is not shared among all pricing networks within $\mathcal{N}_{q}$. Finally, the usage of offset reduced costs (21), jointly with inequality (26), allows for the proof of dominance with respect to the original dominance rule (16)–(20) in all pricing networks in $\mathcal{N}_{q}$ if dominance with respect to (21)–(26) is satisfied in an arbitrary pricing network in $\mathcal{N}_{q}$. For additional details on how the equivalence of pricing networks can be derived using properties i)–iii), we refer to the proof presented in E.